OpenStudy (anonymous):

y = 2tan(theta) . Find the value of tangent, then double it. It's a table with pi/6 through 2pi

jimthompson5910 (jim_thompson5910):

Can you post a screenshot of the full problem?

OpenStudy (anonymous):

yeah, just one moment

OpenStudy (anonymous):

sorry my computer is super slow

jimthompson5910 (jim_thompson5910):

that's fine

jimthompson5910 (jim_thompson5910):

ah, I see now

jimthompson5910 (jim_thompson5910):

so what you do is replace \(\Large \theta\) (greek letter theta) with the numbers in the top row Use the unit circle or a table to find that \[\Large \tan(\theta) = \tan(0) = 0\] \[\Large \tan(\theta) = \tan\left(\frac{\pi}{6}\right) = \frac{\sqrt{3}}{3}\] \[\Large \tan(\theta) = \tan\left(\frac{\pi}{4}\right) = 1\] etc etc

jimthompson5910 (jim_thompson5910):

then you double each result \[\Large 2*\tan(\theta) = \tan(0) = 2*0 = 0\] \[\Large 2*\tan(\theta) = 2*\tan\left(\frac{\pi}{6}\right) = 2*\frac{\sqrt{3}}{3} = \frac{2\sqrt{3}}{3}\] \[\Large 2*\tan(\theta) = 2*\tan\left(\frac{\pi}{4}\right) = 2*1 = 2\] etc etc

OpenStudy (anonymous):

so it's just the coordinate points ?

jimthompson5910 (jim_thompson5910):

yeah the y coordinates of each point on the graph of y = tan(x)

jimthompson5910 (jim_thompson5910):

2*tan(x) I mean

OpenStudy (anonymous):

ugh thank you, you're a life saver

jimthompson5910 (jim_thompson5910):

you're welcome

OpenStudy (anonymous):

wait how'd you get \[\pi/4 \] to be 1

jimthompson5910 (jim_thompson5910):

tan of pi/4 is 1

jimthompson5910 (jim_thompson5910):

if you look at the unit circle, you'll find that sine and cosine have the same value at pi/4 sin(pi/4) = cos(pi/4)

jimthompson5910 (jim_thompson5910):

because of that and because tangent = sine/cosine, the two equal values divide to 1

OpenStudy (anonymous):

oooooooh okay, i get it now

jimthompson5910 (jim_thompson5910):

I'm glad it's making more sense

OpenStudy (anonymous):

Yeah I'm the worst at trigonometry lol

jimthompson5910 (jim_thompson5910):

I'm sure with more practice, you'll get better at it

OpenStudy (anonymous):

this is my third year learning this and I still haven't learned it :/

jimthompson5910 (jim_thompson5910):

then a different approach is needed

OpenStudy (anonymous):

yes, badly so for pi/3 , would it be radical 3 ? or do the 2's not cancel out ?

jimthompson5910 (jim_thompson5910):

the 2s will cancel leaving sqrt(3), correct

OpenStudy (anonymous):

Okay, I just wanted to make sure cause I've seen it where the 2's aren't being canceled out

OpenStudy (anonymous):

and since I'm doubling it, would it be \[2\sqrt{3}\] or ?

jimthompson5910 (jim_thompson5910):

I'd have to see the problem (where you didn't see the 2s cancel), but you should have this \[\Large \tan\left(\theta\right) = \frac{\sin\left(\theta\right)}{\cos\left(\theta\right)}\] \[\Large \tan\left(\frac{\pi}{3}\right) = \frac{\sin\left(\frac{\pi}{3}\right)}{\cos\left(\frac{\pi}{3}\right)}\] \[\Large \tan\left(\frac{\pi}{3}\right) = \frac{\sqrt{3}/2}{1/2}\] \[\Large \tan\left(\frac{\pi}{3}\right) = \frac{\sqrt{3}}{2}\times\frac{2}{1}\] \[\Large \tan\left(\frac{\pi}{3}\right) = \frac{\sqrt{3}}{\cancel{2}}\times\frac{\cancel{2}}{1}\] \[\Large \tan\left(\frac{\pi}{3}\right) = \sqrt{3}\]

jimthompson5910 (jim_thompson5910):

yes, \[\Large 2\tan\left(\frac{\pi}{3}\right) = 2\sqrt{3}\]

OpenStudy (anonymous):

Okay, I think I'm getting a little more. I saw it on some website and it was a table like I'm filling out and the 2's weren't canceled out

jimthompson5910 (jim_thompson5910):

hmm, strange

OpenStudy (anonymous):

but it just makes sense to cancel them out to me so I figured whoever made that table just forgot that step maybe?

jimthompson5910 (jim_thompson5910):

that's possible

jimthompson5910 (jim_thompson5910):

even teachers who write the problems and examples make typos

OpenStudy (anonymous):

I know many lol. But anyways, thank you so much

jimthompson5910 (jim_thompson5910):

sure thing